OpenStudy (anonymous):

Where is the orthocenter of obtuse triangle CDE located? Find GD. A.outside of the triangle; GD = 5 B.inside the triangle; GD = 5 C.outside of the triangle; GD = 7 D.inside the triangle; GD = 7

OpenStudy (mathstudent55):

Is there a figure?

OpenStudy (jacob902):

We know that EGC is the supplementary angle of CGD, so: ∠EGC=180∘−∠CGD=180∘90∘=90∘ But we are also told that: EGC=(5x+55)∘ So we can say: (5x+55)∘=90∘5x+55=905x=90−55=35x=355=7 So: GD=2x−9=2⋅7−9=14−9=5

OpenStudy (jacob902):

Also, for obtuse triangles the orthocenter will be outside the triangle: http://www.mathopenref.com/triangleorthocenter.html

OpenStudy (anonymous):



OpenStudy (jacob902):

almost but still good

OpenStudy (anonymous):